# Planetary Edge Trends (PET)

### v1

@@ 3.1 Dependence of Inner Edge on Stellar Mass
Figure 3: The inner edge of planetary systems as a function of stellar mass for All Multiple Systems, Super-Earth Systems, Sub-Neptune Systems, and Mixed Multiple Systems. The four panels illustrate the correlation between the inner edge and stellar mass for different populations, with different colors used to represent each population: green for All Multiple Systems, red for Super-Earth Systems, blue for Sub-Neptune Systems, and purple for Mixed Multiple Systems. Points represent the observational data, while pentagons indicate the median value for each bin, with error bars showing $1\sigma$. In this work, the binning is merely intended to guide the eye. Dashed lines represent the fit to the observational data based on Eq. (3). The power-law index corresponds to $\gamma_{1}$ in Eqs. (2) and (3), with error bars also indicating $1\sigma$. $r_{obs}$ and $p$ represent the Pearson coefficients and p-values for the four populations, respectively.
We define the inner edge of the planetary system as the location of the innermost planet within the system. Although the orbital period and the semi-major axis are equivalent measures of orbital size, we choose to work with the semi-major axis to study the dependence of the inner edge on stellar mass for better comparison with previous observational and theoretical studies. Thus, the relationship between the inner edge ($a_{\text{in}}$) of planetary systems and stellar mass ($M_{\star}$) can be described by the following simple power-law form:
$a_{\text{in}}=\gamma_{0}M_{\star}^{\gamma_{1}},$ (2)
@@
$\log(a_{\text{in}})=\log(\gamma_{0})+\gamma_{1}\log(M_{\star}).$ (3)
We can calculate $a_{\text{in}}$ by obtaining the planet orbital period (P) from the $Kepler$ catalog, using the following formula:
@@
Using Eq. (3) and applying the Least Squares Method for fitting, we present the observational results of this relationship. From Fig. 3, a correlation between stellar mass and the inner edge can be observed across different planetary populations, with the strength of this correlation ($\gamma_{1}$) varying among them. The four panels represent the four populations mentioned in Sect. 2.3. For all multiple systems, super-Earths, sub-Neptunes, and mixed multiple systems, the correlation index $\gamma_{1}$ is $0.634_{-0.039}^{+0.039}$, $0.543_{-0.052}^{+0.053}$, $0.384_{-0.088}^{+0.087}$, and $0.950_{-0.045}^{+0.046}$, respectively. The uncertainties of these results are calculated by resampling stellar mass within the error range provided in Berger20, recalculating $a_{\text{in}}$ using Eq. (4), and repeating the fitting process 10,000 times. All error bars in this paper represent the $1\sigma$ range.
@@
To further investigate the correlation between stellar mass and the inner edge, we perform a Pearson coefficient test to quantify this correlation. The results are displayed in the bottom right corner of each panel (Fig. 3). It can be observed that all populations exhibit correlations, but the strength and confidence level of these correlations vary. For all multiple systems, super-Earths, sub-Neptunes, and mixed multiple systems, the Pearson coefficients ($r_{obs}$) are 0.211, 0.225, 0.139, and 0.377, respectively. It is worth remarking at this point that these are relatively modest correlations, and we will discuss the scatter in Sect. 4.3. To calculate the p-value, we shuffle the observed data and randomly pair stellar masses and inner edges to create a new dataset matching the size of the observed data. We then calculate the Pearson coefficient ($r$) for this random dataset. Repeating this process 100,000 times ($N_{tot}$), we compare the $r$ values with $r_{obs}$ and count the number of cases where $r>r_{obs}$, denoted as $N(r>r_{obs})$. Thus, the p-value is calculated as $N(r>r_{obs})/N_{tot}$. A smaller p-value indicates a more reliable correlation. The results show that, for all multiple systems, super-Earths, sub-Neptunes, and mixed multiple systems, the p-values ($p$) are 0.0035, 0.0522, 0.2190, and 0.0003, respectively. Therefore, the correlations for all multiple systems and mixed multiple systems are stronger and more significant, while the correlations for super-Earths and sub-Neptunes are weaker, especially for sub-Neptunes. The weaker confidence levels for these two populations may be due to smaller sample sizes. Note that, at this point, we have not considered the influence of other stellar properties. In the following sections, we will further investigate whether this apparent correlation is influenced by other factors.
@@ 3.2 Effect of Stellar Metallicity
Figure 4: Effect of Stellar Metallicity on the Stellar MassInner Edge Correlation. The metallicity dependence of the inner edge and stellar mass, as well as their projection (top two panels) on the inner edge and stellar mass diagram (bottom left panel). Green solid points represent all observational samples, while orange hollow points represent mock datasets generated from the simulation process. The light green dashed line represents the fitting results for the green solid points ($\gamma_{2}$ and $\gamma_{3}$ in Eqs. (5) and (6)), while the orange dashed line represents the fitting results for the orange hollow points ($\gamma_{1}$ in Eq. (3)). The top two panels and the bottom left panel collectively display the median result from the 10,000 simulation runs. The bottom right panel shows the distribution of the correlation index ($\gamma_{1}$) between the inner edge and stellar mass for the mock data (using Kernel Density Estimation), with the orange dashed line indicating the median of the distribution, the black dashed line indicating a correlation of zero, and the green dashed line representing the correlation of all observational data (as shown in the top left panel of Fig. 3). Note: To further improve our analysis of stellar metallicity, we remove two outliers (with values less than $-1.0$) from the correlation analysis.
Since stellar metallicity influences planetary architecture (e.g., orbital periods) and is also correlated with stellar mass, the effect of metallicity should be considered in order to remove its impact on the relationship between the inner edge and stellar mass. We plot the top two panels of Fig. 4, showing the inner edge as a function of metallicity and stellar mass as a function of metallicity, respectively. As can be seen from the fitting results of the observed data (represented by the green solid points and the light green dashed lines), the inner edge exhibits a negative correlation with metallicity ($\gamma_{2}=-0.201$), while metallicity shows a positive correlation with stellar mass ($\gamma_{3}=0.094$). To investigate the projection of these two relationships on the inner edgestellar mass diagram, we conduct the following data mocking and fitting analysis.
We generate the mock data of the inner edge and stellar mass, with the specific calculations corresponding to the following functional forms:
$\log(a_{\text{in}}^{\text{mock}})=\log(\gamma_{0})+\gamma_{2}\,\text{[Fe/H]}+\text{res}(a_{\text{in}}),$ (5)
@@
$\log(M_{\star}^{\text{mock}})=\log(\gamma_{0})+\gamma_{3}\,\text{[Fe/H]}+\text{res}(M_{\star}),$ (6)
where $\gamma_{0}$ and $\gamma_{2}$ in Eq. (5) correspond to the best fit in the top left panel of Fig. 4, and $\gamma_{0}$ and $\gamma_{3}$ in Eq. (6) correspond to the best fit in the top right panel of Fig. 4. $\text{res}(a_{\text{in}})$ and $\text{res}(M_{\star})$ are the residuals, which are randomly selected from the residuals of the top left panel and top right panel, respectively. As shown in the bottom left panel of Fig. 4, we show a typical case of such a data mocking and fitting analysis. Fitting this mock data (orange hollow points) using Eq. (3) produced an anti-correlation on the $a_{\text{in}}-M_{\star}$ diagram (represented by the orange dashed line with a $\gamma_{1}$ power-law index value of $-0.128$).
The above data mocking and fitting process is repeated 10,000 times, and the distribution of the correlation index ($\gamma_{1}$) between the inner edge and stellar mass for the mock data is illustrated in the bottom right panel. As can be seen, the median power-law index of the correlation distribution is $-0.128$. In contrast, our observational results show a positive correlation between the inner edge and stellar mass of 0.634 (the correlation for all multiple systems, as shown in the top left panel of Fig. 3), which is significantly larger than the median of the distribution. Furthermore, out of 10,000 data mocking and fitting cases, only 7 produced a positive correlation with $\gamma_{1}>0.634$, corresponding to a p-value of 0.0007 for a hypothesis that the observed positive correlation is caused by the projection of the above metallicity effect (Eqs. (5) and (6)).
@@
Figure 5: The dependence of the inner edge on stellar mass and metallicity, as well as the corresponding dependence between mass and metallicity, obtained through the MLR model. The left and middle panels show the power-law indices, which correspond to $\gamma_{1}$ and $\gamma_{2}$ in Eqs. (7) and (8), respectively. Additionally, the right panels show the power-law index corresponding to $\gamma_{3}$ in Eq. (9). The lines represent the fitting results, with color bands indicating $1\sigma$ error bars. The colors and styles of points and lines are consistent with those in Fig. 3.
@@ 3.3 Multiple Linear Regression Model
To investigate the dependence of the inner edge on both stellar mass and metallicity, we utilize the Multiple Linear Regression (MLR) model. MLR identifies the best-fitting straight line to describe the relationship between multiple input variables and the output variable. Although the relationship between stellar mass, metallicity, and the inner edge is not strictly linear, we can convert it into a multiple linear relationship through logarithmic transformation and analyze it using a multiple linear regression model. To better express the relationships between stellar mass, metallicity, and the inner edge, we adopt the following relational equation, a simple power-law model in the form:
$\frac{a_{\text{in}}}{\text{AU}}=\gamma_{0}\left(\frac{M_{\star}}{M_{\odot}}\right)^{\gamma_{1}}10^{\gamma_{2}\,\text{[Fe/H]}}.$ (7)
The equation becomes easier to work with in logarithmic space:
@@
Using the MLR model from Eq. (8), we derive the dependence of the inner edge on stellar mass and metallicity for each population dataset. In Fig. 5, the four rows of panels present the results of the four populations obtained through the MLR model. The left panels show the distribution of the inner edge and stellar mass, as well as the fitting results for all multiple systems, super-Earths, sub-Neptunes, and mixed multiple systems, with $\gamma_{1}$ values of $0.812_{-0.078}^{+0.085}$, $0.566_{-0.108}^{+0.100}$, $0.667_{-0.182}^{+0.173}$, and $1.115_{-0.072}^{+0.082}$, respectively. The middle panels display the distribution of the inner edge and metallicity, as well as the fitting results, with $\gamma_{2}$ values of $-0.277_{-0.026}^{+0.028}$, $-0.179_{-0.043}^{+0.035}$, $-0.182_{-0.070}^{+0.105}$, and $-0.286_{-0.031}^{+0.029}$, respectively. Additionally, we determine the corresponding relationship between stellar mass and metallicity by fitting it using the Least Squares Method through the following equation:
@@
The right panels in Fig. 5 show the distribution of stellar mass and metallicity, as well as the fitting results, with $\gamma_{3}$ values of $0.094_{-0.010}^{+0.011}$, $0.019_{-0.017}^{+0.017}$, $0.169_{-0.018}^{+0.018}$, and $0.093_{-0.018}^{+0.018}$, respectively.
According to Fig. 5, we observe that the inner edge shows a positive correlation with mass and a negative correlation with metallicity. From all multiple systems and mixed multiple systems populations, we observe that, under similar influences of metallicity on the inner edge (middle panels showing $\gamma_{2}$ values of approximately $-0.277$ and $-0.286$, respectively), the correlation between the inner edge and stellar mass increases by a similar magnitude after correcting for the metallicity effect (left panels compared to the results in Fig. 3, $\gamma_{1}$ values of approximately 0.2). However, in super-Earths and sub-Neptunes populations, under similar influences of metallicity on the inner edge (middle panels showing $\gamma_{2}$ values of approximately $-0.179$ and $-0.182$, respectively), the correlation between the inner edge and stellar mass shows varying degrees of increase after correcting for the metallicity effect. For super-Earths, the increase is minimal (left panel compared to the result in Fig. 3, $\gamma_{1}$ value of approximately 0.02), whereas for sub-Neptunes, the increase is more pronounced (left panel compared to the result in Fig. 3, $\gamma_{1}$ value of approximately 0.3). We believe this difference is primarily due to the correlation between stellar mass and metallicity ($\gamma_{3}$ in Eq. (9)). The values of $\gamma_{3}$ for all multiple systems and mixed multiple systems are similar (right panels showing $\gamma_{3}$ values of approximately 0.094 and 0.093, respectively), but there is a significant difference between super-Earths and sub-Neptunes (right panels showing $\gamma_{3}$ values of approximately 0.019 and 0.169, nearly an order of magnitude). Sub-Neptunes have the largest $\gamma_{3}$, while super-Earths have the smallest $\gamma_{3}$. Therefore, the middle and right panels together determine the degree of metallicity correction, and the strength of the correction effect varies across the four populations. This may explain the varying degrees of increase (the change in $\gamma_{1}$ from Fig. 3 to Fig. 5) observed among the four populations.

### v2

@@ 3.1 Dependence of inner edge on stellar mass
Figure 3: The inner edge of planetary systems as a function of stellar mass for the datasets of all multiple systems, super-Earth systems, sub-Neptune systems, and mixed multiple systems. The four panels illustrate the correlation between the inner edge and stellar mass for different populations. Different colors are used to represent each population: green for all multiple systems, red for super-Earth systems, blue for sub-Neptune systems, and purple for mixed multiple systems. Points represent the observational data, while pentagons indicate the median value for each bin, with error bars showing $1\sigma$. In this work, the binning is merely intended to guide the eye. Dashed lines represent the fit to the observational data based on Eq. (3). The power-law index corresponds to $\gamma_{1}$ in Eqs. (2) and (3), with error bars also indicating $1\sigma$. $r_{obs}$ and $p$ represent the Pearson coefficients and p-values for the four populations, respectively. Note: The median values and $1\sigma$ dispersion of the binned data in this study are calculated in logarithmic space and are based on actual observational data.
In this study, we primarily focused on the relationship between stellar mass and the inner edge. For stellar effective temperature (Pecaut & Mamajek, 2013), we selected main-sequence stars, where temperature and mass are approximately equivalent. An empirical relation exists between them, given by $T_{\text{eff}}\propto M_{\star}^{0.5}$. Therefore, by studying stellar mass, we also indirectly account for stellar temperature. Stellar age is another potentially important factor; however, we did not include it in this work due to the large uncertainties in current age estimates (Berger et al., 2020).
We defined the inner edge of the planetary system as the location of the innermost planet within the system. Although the orbital period and the semimajor axis are equivalent measures of orbital size, we chose to work with the semimajor axis to study the dependence of the inner edge on stellar mass for better comparison with previous observational and theoretical studies. Thus, the relationship between the inner edge ($a_{\text{in}}$) of planetary systems and stellar mass ($M_{\star}$) can be described by the following simple power-law form:
$a_{\text{in}}={\color[rgb]{0,0,0}\gamma_{0,am}}M_{\star}^{\gamma_{1}},$ (2)
@@
$\log(a_{\text{in}})={\color[rgb]{0,0,0}\log(\gamma_{0,am})}+\gamma_{1}\log(M_{\star}).$ (3)
It is important to note that in this study, all $\gamma_{0}$ terms refer to the intercept, with different subscript letters indicating the specific combination of parameters included in each equation. We can calculate $a_{\text{in}}$ by obtaining the planet orbital period (P) from the $Kepler$ catalog using the following formula:
@@
Figure 4: Effect of stellar metallicity on the stellar massinner edge correlation. The metallicity dependence of the inner edge and stellar mass, as well as their projection (top two panels) on the inner edge and stellar mass diagram (bottom-left panel). Green solid points represent all observational samples, while orange hollow points represent mock datasets generated from the simulation process. The light green dashed line represents the fitting results for the green solid points ($\gamma_{2}$ and $\gamma_{3}$ in Eqs. (5) and (6)), while the orange dashed line represents the fitting results for the orange hollow points ($\gamma_{1}$ in Eq. (3)). The top two panels and the bottom-left panel collectively display the median result from the 10,000 simulation runs. The bottom-right panel shows the distribution of the correlation index ($\gamma_{1}$) between the inner edge and stellar mass for the mock data (using Kernel Density Estimation), with the orange dashed line indicating the median of the distribution, the black dashed line indicating a correlation of zero, and the green dashed line representing the correlation of all observational data (as shown in the top-left panel of Fig. 3). Note: To further improve our analysis of stellar metallicity, we removed two outliers (with values less than $-1.0$) from the correlation analysis.
@@
Using Eq. (3) and applying the least squares method for fitting, we presented the observational results of this relationship. From Fig. 3, a correlation between stellar mass and the inner edge can be observed across different planetary populations, with the strength of this correlation ($\gamma_{1}$) varying among them. The four panels represent the four populations mentioned in Sect. 2.3. For all multiple systems, super-Earths, sub-Neptunes, and mixed multiple systems, the correlation index $\gamma_{1}$ is $0.63_{-0.04}^{+0.04}$, $0.54_{-0.05}^{+0.05}$, $0.38_{-0.09}^{+0.09}$, and $0.95_{-0.05}^{+0.05}$, respectively. The uncertainties of these results were calculated by resampling stellar mass within the error range provided in Berger20, recalculating $a_{\text{in}}$ using Eq. (4), and repeating the fitting process 10,000 times. All error bars in this paper represent the $1\sigma$ range.
@@
To further investigate the correlation between stellar mass and the inner edge, we performed a Pearson coefficient test to quantify this correlation. The results are displayed in the bottom-right corner of each panel (Fig. 3). It can be observed that all populations exhibit correlations, but the strength and confidence level of these correlations vary. For all multiple systems, super-Earths, sub-Neptunes, and mixed multiple systems, the Pearson coefficients ($r_{obs}$) are 0.211, 0.225, 0.139, and 0.377, respectively. It is worth remarking at this point that these are relatively modest correlations, and we discuss the scatter in Sect. 4.3. To calculate the p-value, we shuffled the observed data and randomly paired stellar masses and inner edges to create a new dataset matching the size of the observed data. We then calculated the Pearson coefficient ($r$) for this random dataset. Repeating this process 100,000 times ($N_{tot}$), we compared the $r$ values with $r_{obs}$ and counted the number of cases where $r>r_{obs}$, denoted as $N(r>r_{obs})$. Thus, the p-value was calculated as $N(r>r_{obs})/N_{tot}$. A smaller p-value indicates a more reliable correlation. The results show that, for all multiple systems, super-Earths, sub-Neptunes, and mixed multiple systems, the p-values ($p$) are 0.0032, 0.0506, 0.2177, and 0.0003, respectively. Therefore, the correlations for all multiple systems and mixed multiple systems are stronger and more significant, while the correlations for super-Earths and sub-Neptunes are weaker, especially for sub-Neptunes. The weaker confidence levels for these two populations may be due to smaller sample sizes. Here, we performed simulations to test the quantitative impact of sample size on the results for super-Earths and sub-Neptunes. This was tested by randomly drawing the same number of planets from each population within the all multiple systems sample and analyzing the resulting distribution of slope values over 10,000 simulations. The results show that the median and $1\sigma$ range of the slope distributions for super-Earths and sub-Neptunes are $0.63_{-0.35}^{+0.35}$ and $0.65_{-0.49}^{+0.49}$, respectively. It is clear that both distributions have large error bars, and the observed slopes of $0.54_{-0.05}^{+0.05}$ and $0.38_{-0.09}^{+0.09}$ fall within the $1\sigma$ range of the simulated distributions. Therefore, we cannot rule out the possibility that the observed difference between super-Earths and sub-Neptunes is due to limited sample sizes.
We note that, at this point, we have not considered the influence of other stellar properties. In the following sections, we further investigate whether this apparent correlation is influenced by other factors.
@@ 3.2 Effect of stellar metallicity
Previous studies have demonstrated that stellar metallicity influences the architecture of planetary systems. Beaugé & Nesvorný (2013) reported, based on statistical analysis, that stellar metallicity affects the distribution of planetary periods, which could, in turn, influence the position of the inner edge. Similarly, Mulders et al. (2016) found that short-period planets—particularly hot rocky planets—are more likely to orbit host stars with higher metallicity.
Since stellar metallicity influences planetary architecture (e.g., orbital periods) and is also correlated with stellar mass, the effect of metallicity should be considered in order to remove its impact on the relationship between the inner edge and stellar mass. We plot the top two panels of Fig. 4, showing the inner edge as a function of metallicity and stellar mass as a function of metallicity, respectively. As can be seen from the fitting results of the observed data (represented by the green solid points and the light green dashed lines), the inner edge exhibits a negative correlation with metallicity ($\gamma_{2}=-0.20$), while metallicity shows a positive correlation with stellar mass ($\gamma_{3}=0.09$). To investigate the projection of these two relationships on the inner edgestellar mass diagram, we conducted the following analysis of data mocking and fitting.
We generated the mock data of the inner edge and stellar mass with the specific calculations corresponding to the following functional forms:
$\log(a_{\text{in}}^{\text{mock}})={\color[rgb]{0,0,0}\log(\gamma_{0,afr})}+\gamma_{2}\,\text{[Fe/H]}+\text{res}(a_{\text{in}})$ (5)
@@
$\log(M_{\star}^{\text{mock}})={\color[rgb]{0,0,0}\log(\gamma_{0,mfr})}+\gamma_{3}\,\text{[Fe/H]}+\text{res}(M_{\star}),$ (6)
where $\gamma_{0,afr}$ and $\gamma_{2}$ in Eq. (5) correspond to the best fit in the top-left panel of Fig. 4, and $\gamma_{0,mfr}$ and $\gamma_{3}$ in Eq. (6) correspond to the best fit in the top-right panel of Fig. 4. The terms $\text{res}(a_{\text{in}})$ and $\text{res}(M_{\star})$ are the residuals, which are randomly selected from the residuals of the top-left panel and top-right panel, respectively. In the bottom-left panel of Fig. 4, we show a typical case of data mocking and fitting. Fitting this mock data (orange hollow points) using Eq. (3) produced an anticorrelation on the $a_{\text{in}}-M_{\star}$ diagram (represented by the orange dashed line with a $\gamma_{1}$ power-law index value of $-0.13$).
The above data mocking and fitting process was repeated 10,000 times, and the distribution of the correlation index ($\gamma_{1}$) between the inner edge and stellar mass for the mock data is illustrated in the bottom-right panel of Fig. 4. As can be seen, the median power-law index of the correlation distribution is $-0.13$. In contrast, our observational results show a positive correlation between the inner edge and stellar mass of 0.63 (the correlation for all multiple systems, as shown in the top-left panel of Fig. 3), which is significantly larger than the median of the distribution. Furthermore, out of 10,000 data mocking and fitting cases, only six produced a positive correlation with $\gamma_{1}>0.63$, corresponding to a p-value of 0.0006 for a hypothesis that the observed positive correlation was caused by the projection of the above metallicity effect (Eqs. (5) and (6)).
@@ 3.3 Multiple linear regression model
Figure 5: Dependence of the inner edge on stellar mass and metallicity as well as the corresponding dependence between mass and metallicity obtained through the MLR model. The left and middle panels show the power-law indices, which correspond to $\gamma_{1}$ and $\gamma_{2}$ in Eqs. (7) and (8), respectively. Additionally, the right panels show the power-law index corresponding to $\gamma_{3}$ in Eq. (9). The lines represent the fitting results, with color bands indicating $1\sigma$ error bars. The colors and styles of points and lines are consistent with those in Fig. 3. Note: The median values and $1\sigma$ dispersion of the binned data in this study are calculated in logarithmic space and are based on actual observational data.
To investigate the dependence of the inner edge on both stellar mass and metallicity, we utilized the multiple linear regression (MLR) model. The MLR identifies the best-fitting straight line to describe the relationship between multiple input variables and the output variable. Although the relationship between stellar mass, metallicity, and the inner edge was not strictly linear, we converted it into a multiple linear relationship through logarithmic transformation and analyzed it using an MLR model. To better express the relationships between stellar mass, metallicity, and the inner edge, we adopted the following relational equation, a simple power-law model in the form:
$\frac{a_{\text{in}}}{\text{AU}}={\color[rgb]{0,0,0}\gamma_{0,amf}}\left(\frac{M_{\star}}{M_{\odot}}\right)^{\gamma_{1}}10^{\gamma_{2}\,\text{[Fe/H]}}.$ (7)
The equation became easier to work with in logarithmic space:
@@
Equation (7) mainly follows previous studies on occurrence rates, such as Eq. (1) in Johnson et al. (2010) and Eq. (21) in Zhu & Dong (2021). However, these formulas do not have a strong theoretical or astrophysical basis. Although they are not specifically designed for studying the inner edge, we believe they can serve as a starting point for both preliminary and detailed analyses. More importantly, using these formula forms allows us to directly compare our results with previous observational findings and theoretical models (see Sect. 4), thereby making the results more meaningful for interpretation and discussion.
Using the MLR model from Eq. (8), we derived the dependence of the inner edge on stellar mass and metallicity for each population dataset. We adopted the same method as in Sect. 3.1 to report the uncertainties for the slopes, which were derived by accounting for individual data point uncertainties and repeating the process 10,000 times. In Fig. 5, the four rows of panels present the results of the four populations obtained through the MLR model. The left panels show the distribution of the inner edge and stellar mass, as well as the fitting results for all multiple systems, super-Earths, sub-Neptunes, and mixed multiple systems, with $\gamma_{1}$ values of $0.81_{-0.08}^{+0.09}$, $0.57_{-0.11}^{+0.10}$, $0.67_{-0.18}^{+0.17}$, and $1.12_{-0.07}^{+0.08}$, respectively. The middle panels display the distribution of the inner edge and metallicity, as well as the fitting results, with $\gamma_{2}$ values of $-0.28_{-0.03}^{+0.03}$, $-0.18_{-0.04}^{+0.04}$, $-0.18_{-0.07}^{+0.11}$, and $-0.29_{-0.03}^{+0.03}$, respectively. Additionally, we determined the corresponding relationship between stellar mass and metallicity by fitting it using the least squares method through the following equation:
@@
The right panels in Fig. 5 show the distribution of stellar mass and metallicity, as well as the fitting results, with $\gamma_{3}$ values of $0.09_{-0.01}^{+0.01}$, $0.02_{-0.02}^{+0.02}$, $0.17_{-0.02}^{+0.02}$, and $0.09_{-0.02}^{+0.02}$, respectively.
According to Fig. 5, we observe that the inner edge shows a positive correlation with mass and a negative correlation with metallicity. From all multiple systems and mixed multiple systems populations, we observe that, under similar influences of metallicity on the inner edge (middle panels showing $\gamma_{2}$ values of approximately $-0.28$ and $-0.29$, respectively), the correlation between the inner edge and stellar mass increases by a similar magnitude after correcting for the metallicity effect (left panels compared to the results in Fig. 3, $\gamma_{1}$ values of approximately 0.2). However, in super-Earths and sub-Neptunes populations, under similar influences of metallicity on the inner edge (middle panels showing $\gamma_{2}$ values of approximately $-0.18$), the correlation between the inner edge and stellar mass shows varying degrees of increase after correcting for the metallicity effect. For super-Earths, the increase is minimal (left panel compared to the result in Fig. 3, $\gamma_{1}$ value of approximately 0.03), whereas for sub-Neptunes, the increase is more pronounced (left panel compared to the result in Fig. 3, $\gamma_{1}$ value of approximately 0.3). We believe this difference is primarily due to the correlation between stellar mass and metallicity ($\gamma_{3}$ in Eq. (9)). The values of $\gamma_{3}$ for all multiple systems and mixed multiple systems are similar (right panels showing $\gamma_{3}$ values of approximately 0.09), but there is a significant difference between super-Earths and sub-Neptunes (right panels showing $\gamma_{3}$ values of approximately 0.02 and 0.17, nearly an order of magnitude). Sub-Neptunes have the largest $\gamma_{3}$, while super-Earths have the smallest $\gamma_{3}$. Therefore, the middle and right panels together determine the degree of metallicity correction, and the strength of the correction effect varies across the four populations. This may explain the varying degrees of increase (the change in $\gamma_{1}$ from Fig. 3 to Fig. 5) observed among the four populations.
@@
We note that modifying the functional form may lead to different strengths of the positive correlation between the inner edge and stellar mass, as well as the negative correlation between the inner edge and stellar metallicity. After testing various alternatives, we conclude that caution should be exercised when interpreting the exact numerical values. However, given the current state of research, the functional form adopted in this study is considered reasonable, as it helps to explain the observed trends and facilitates direct comparison with previous observational studies and theoretical models (see Sect. 4), thereby leading to more meaningful results and conclusions.